West High School's population is 250 studens less than twice the population of East high school. combined they serve 2,858 students. How many students attend West High School: West High School's population is 250 studens less than twice the population of East high school. combined they serve 2,858 students. How many students attend West High School
Answer by orca(336)

Let x be the number of students attending East High School.
The the number of students attending West High School can be expressed in terms of x as:
2x - 250
The total number of students in the two schools is x + 2x - 250
As the total is 2858, we have
x + 2x - 250 = 2858
Solving for x, we have
3x - 250 = 2858
3x = 3108
x = 1036
So the number of students attending East High School is 1036.
The number of students attending West High School can be found by substituting x = 1036 into 2x - 250.
2x - 250 = 2*1036 - 250 = 1822